OpenStudy (anonymous):

Solve the following log without using a calculator. 10^x^2-4 = 1

OpenStudy (anonymous):

\[10^{x^2-4}=1\]

OpenStudy (anonymous):

So far using the prop of logs, I have: \[\log_{10}10^{x^2-4}=\log_{10}1 \]

OpenStudy (anonymous):

canceling that log10 on the left gives me: \[x^2-4\] but I'm not sure what I do for the right side.

OpenStudy (anonymous):

\[ 10^{0}=1 \]

OpenStudy (yttrium):

You can use x^2 - 4 = 0 Then solve x

OpenStudy (akashdeepdeb):

In Mathematics you should NEVER Canel logs.

OpenStudy (anonymous):

@wio Oh right, I forgot about that.

OpenStudy (anonymous):

From there I just move the 4 over to the left side and square both sides to get the following: \[x=\pm \sqrt{2}\] right?

OpenStudy (anonymous):

how did you get \(\sqrt 2\)?

OpenStudy (anonymous):

oh sorry it would just be -2 and +2 Could I also factor it to get: \[(x+2)(x-2)=0\] \[x=\pm2\]